# Diffusing in Someone Else’s Shoes: Robotic Perspective Taking with Diffusion

### v1

@@ IV Results
Figure 3: One example of predicting the first-person view, where the direct prediction differs from the prediction after iterative denoising. In the direct prediction, two thumbs can be identified on the hand, whereas only one is left after the iterative denoising, although in the wrong position.
Figure 4: Qualitative comparisons between our model and pix2pix as well as CycleGAN. Five examples (A,B,C,D,E) are shown to illustrate the differences between the models. CycleGan has almost no variance between its output, pix2pix mostly seems to have the correct poses except for example C, but often is unable to correctly recreate the hands and finger structure. Our model always has the correct pose and recreates the hands in a lower resolution.
We ran multiple versions of our model to evaluate the strengths and weaknesses. First off, we have the image-to-image model that transfers an image from the third-person perspective to the first-person perspective. Both qualitative and quantitative evaluations were performed.
We compare the image generation part of our approach to the pix2pix [8] model and the CycleGAN [9] model. We make use of the implementation provided by the authors of these papers [^3] https://github.com/junyanz/pytorch-CycleGAN-and-pix2pix/tree/master and use the standard configurations for both models to train them on our dataset. We evaluate both of these models as well as our own on three metrics with our test set. The metrics used are mean square error (MSE), L1 norm (L1) and the structural similarity index measure (SSIM) [27]. For MSE and L1, smaller is better, while the opposite is true for the SSIM. Our model outperformed both comparison models across all three metrics, as shown in Tab. I.
Figure 5: Joint predictions from third-person image.
@@
Model MSE L1 SSIM
@@
Qualitative analysis, Fig. 4, shows that CycleGan seems to be unable to correctly adjust to the condition, always generating the same pose for the first-person perspective. Pix2pix goes one step further with the generated image showing an understanding of the translation from third- to first-person. Compared to our result, however, the hands are often impossible to make out; there are still significant differences in the position and size of the arms and in some cases, artefacts appear in the image. Our results only seem to differ in the resolution compared to the goal images taken directly from the simulation. The hands and arms are in the correct positions, with even the fingers being clearly visible and discernible.
@@
For the iterative denoising, we tried out doing 0 steps, 50 steps, 200 steps or the full 1000 steps. In most of our results, the differences between the goal and the image resulting either from a direct prediction or from a prediction after going through any number of iterative denoising steps are seemingly impossible to make out, with the prediction from the first step already looking very close to the ground truth. We did, however, come across one image, as seen in Fig. 3, in which finding the right position of the thumb was difficult for the model, resulting in two thumbs for the direct prediction, which was reduced to one thumb after the denoising process although in the wrong position. There are also cases where some of the blur from the goal images captured in the simulation was smoothed out in the predictions.
When it comes to predicting the joints, we have again shown both quantitative, through the mean square error on our validation set, and qualitative results, by showing images of the poses resulting from the predicted joints. Our average mean squared error over the validation set is 27e-4 for the joints. It appears that the model finds it challenging to generalise from the training set, where we are able to get to a loss as low as 3.3e-6. While the general configuration seems to go in the correct direction, often, some of the joints are not completely accurate, shifting the whole arm in a wrong direction. While not as good as the image-generating part, the model still demonstrated an ability to infer a relationship between the condition and the goal joint configuration as seen in Fig. 5.
We combined the results introduced so far, by training our joint position model on the first-person images from the demonstration. As the first-person images are supposed to be easier to learn from, and we are capable of generating them, it makes sense to see if learning the joint values from them is easier. By changing the input images, we managed to reduce the training error to 3e-7 and the validation error to 0.0017, clearly showing the use of the first-person perspective. Augmenting the input data with rotations, flips or maskings was able to further improve the validation error to 0.0014 where we reached a limit. One example of the results from using first-person images for the condition is shown in Fig. 6.
It seems to be more difficult to directly predict the joints instead of recreating the first-person image, despite both requiring a spatial understanding through the third-person image. One possible reason for that is likely the redundancy in the arm. Allowing for many poses that look similar while having very different joint configurations. Fig. 6 shows one particular interesting result demonstrating this, where the first-person view looks very similar, while the actual pose looks quite different. With so many very similar looking poses, the model tends to predict values closer to the mean.
To evaluate the predicted joints in more detail, Tab. II shows the mean square error for each of the joints in the arm; we did not evaluate the five-finger joint values. All versions of our model have an error far below the standard deviation of the dataset. The individual errors also show a very close relation to the existing standard deviation. Even though different parts of the robot are visible in the first-person view compared to the third-person view the different models do not show a large discrepancy between the visible and non-visible parts.
@@
Figure 6: One example of predicting joint values from the first-person view, where the imitation looks similar from the first-person view while looking very different from the third-person view.
TABLE II: The mean squared error for each of the joint values. All the values are e-4. The prediction error is correlating with the standard deviation of the given joint.
Model j1 j2 j3 j4 j5 j6 j7 j8
Left arm
dataset std 297 762 914 180 436 153 266 294
third-person 14 79 111 4 28 3 11 13
first-person 12 71 95 4 24 3 10 10
augmented 11 67 82 3 20 3 7 9
Right arm
dataset std 271 723 860 114 307 119 279 390
third-person 11 69 96 2 14 2 12 22
first-person 10 61 85 2 13 2 10 19
augmented 16 52 78 2 10 2 18 16
@@ V Conclusion
We have presented that we can generate images in first-person perspective from a third-person perspective that are close to indistinguishable from the real first-person perspective, surpassing established approaches such as pix2pix or CycleGAN. While significant progress has been made, challenges remain in imitation learning from a third-person perspective. We believe that adding a pose detector to our model to get a more direct representation of the demonstrated pose can be a fruitful improvement, especially when extending the approach to the real world. Adding more three-dimensional information like this and making use of models focused on detecting arm poses should make it easier for our model to understand which specific joint configuration is needed. Another way of continuing this work would be to add an already existing approach that uses image data from an ego perspective to imitate poses.
Throughout this paper, we have demonstrated that our approach effectively learns spatial relations and can resize and rotate objects to adjust them to a new perspective. Our results also highlight the utility of first-person images in learning joint positions, which significantly reduced our loss compared to third-person images. Additionally, we have demonstrated the potential of diffusion models for pose estimation and third-person imitation.

### v2

@@ IV Results
In this section, we first evaluate our architecture by comparing it with other image-to-image models. The robustness of our model is tested by evaluating it on additional datasets. Finally, we look at how perspective-taking can improve the generation of joint values.
IV-A Comparison with pix2pix and CycleGan
Figure 4: One example of predicting the first-person perspective, where the direct prediction differs from the prediction after the iterative denoising. In the direct prediction, two thumbs can be identified on the hand, whereas only one is left after the iterative denoising, although in the wrong position. The resolution is in accordance with the model’s input and output.
We compare the image generation part of our approach to the pix2pix [10] model and the CycleGAN [11] model. We use the implementation provided by the authors of these papers [^2] https://github.com/junyanz/pytorch-CycleGAN-and-pix2pix/tree/master and use the standard configurations for both models to train them on our dataset. We evaluate both of these models as well as our own on three metrics with our test set. The metrics used are mean square error (MSE), L1 norm (L1) and the structural similarity index measure (SSIM) [27]. MSE and L1 simply measure the difference in pixels, while the SSIM uses luminance, contrast and structural comparisons. For MSE and L1, smaller is better, while the opposite is true for the SSIM. Our model outperformed both comparison models across all three metrics, as shown in Tab. I.
@@ IV-A Comparison with pix2pix and CycleGan
Model MSE $\downarrow$ L1 $\downarrow$ SSIM $\uparrow$
@@
Qualitative analysis shows that CycleGAN is unable to correctly adjust to the condition, consistently generating the same image for the first-person perspective, as illustrated in Fig. 3. Pix2pix goes one step further with the generated image showing an understanding of the translation from third- to first-person. Compared to our result, however, the hands are often impossible to make out; there are still significant differences in the position and size of the arms and in some cases, artefacts appear in the image. Our results only seem to differ in the resolution compared to the goal images taken directly from the simulation. The hands and arms are in the correct positions, with even the fingers being clearly visible and discernible.
@@
Only utilising a single iterative denoising step led to the same MSE of 0.0007, L1 of 0.0086 and a very similar SSIM of 0.9793 showing that the direct prediction of our model can compete with the iteratively denoised prediction. While the quantitative results showed next to no differences from the iterative denoising, we did come across a few examples where the results changed through iterative denoising. In Fig. 4, one example is shown, in which finding the right position of the thumb was difficult for the model, resulting in two thumbs for the direct prediction, which was reduced to one thumb after the denoising process although in the wrong position. There are also cases where some of the blur from the goal images captured in the simulation was smoothed out in the predictions.
IV-B Comparison with EMIL and RH20T
Figure 5: Third-person perspective generation for the EMIL dataset. The resolution is in accordance with the model’s input and output.
Figure 6: Third-Person perspective generation for the RH20T dataset. The resolution is in accordance with the model’s input and output.
Figure 7: Comparison for the development over the iterative denoising for each of the datasets, shown are the generated images after step 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000 from left to right.
To further evaluate the model, we test it on two additional datasets: the EMIL dataset [3] and the RH20T dataset [4], again the model was trained on 80% of these datasets and tested on the remaining 20%. While our dataset was collected in a simulated environment, both of these datasets were collected in the real world. Compared to the simulated environment, there are far more details in the collected frames of the real world. In the EMIL dataset, the background is mostly kept clean, while the RH20T dataset has various shapes and colour patterns in each frame. The RH20T dataset also has the added challenge of changing table covers, while the only thing that changes in the EMIL dataset is the small object, and no external objects change in the simulated dataset. The datasets have different levels of complexity that might be encountered in the field of imitation learning.
For each dataset, we trained a model with the exact same architecture and used the SSIM, MSE and L1 metrics to compare them. We also tested different amounts of iterative denoising steps, as shown in Tab.II, to see if the complexity of the dataset had an influence on the behaviour during inference. The results align with our expectations, indicating that the simulated dataset is the least challenging followed by the EMIL dataset and finally the RH20T dataset. Even for the more complex datasets, the model is capable of realistically generating new perspectives. In the EMIL dataset, the difference between the real first-person perspective and the generated first-person perspective is difficult to discern, even without iterative denoising steps, as shown in Fig. 5. For the RH20T dataset, the model is also able to generate a new perspective, although small details remain challenging. There is a significant difference between the generation without any iterative denoising steps and the generation after 50 iterative denoising steps, as shown in Fig. 6.
The more difficult datasets necessitate iterative denoising to generate finer details and to de-blur parts of the image. Surprisingly, this is not reflected in the quantitative results, likely because the differences only affect small parts of the images. Furthermore, sharper images do not necessarily result in better pixel matchings, as they might be slightly misaligned. To further explore the effect of the iterative denoising, we show one example for each dataset in Fig. 7, in which a full 1000 iterative denoising steps are used during the inference, of which every one-hundredth step is displayed. Only the last jump in denoising steps shows a change. For the RH20T dataset, the robot arm becomes much clearer, with the right-most part only appearing during the last step. In the EMIL dataset, the last step makes the fingertips of NICO’s right hand clearer.
@@ IV-B Comparison with EMIL and RH20T
Figure 8: One example of predicting joint values from the first-person perspective, where the imitation looks similar in the first-person perspective but different in the third-person perspective compared to the demonstration.
TABLE II: Comparison between datasets and amount of iterative denoising steps. Errors are consistent across datasets, with higher errors for the more complex datasets.
Model MSE $\downarrow$ L1 $\downarrow$ SSIM $\uparrow$
Simulated (1 steps) 0.0007 0.0086 0.9793
EMIL (1 steps 0.0014 0.0183 0.9151
RH20T(1 steps 0.0101 0.0603 0.7606
Simulated (200 steps) 0.0007 0.0083 0.9805
EMIL (200 steps 0.0015 0.0181 0.9156
RH20T(200 steps 0.0133 0.0610 0.7628
IV-C Joint Generation
To assert the benefit of the generated first-person perspective, we adjust our architecture by changing the last layer to output joint values instead of RGB images. Then, we train the adjusted architecture with the third-person images once and once with the generated first-person images. We use the simulated NICOL dataset for this purpose. Using 20% of the dataset for validation, the model achieves a mean squared error of 0.0027 for the joint values with the third-person images. It appears that the model finds it challenging to generalise from the training set, where we are able to get to a loss as low as 3.3e-6. While the general configuration seems to go in the correct direction, often, some of the joints are not completely accurate, shifting the whole arm in a wrong direction. Training on the generated first-person images, instead we managed to reduce the training error to 3e-7 and the validation error to 0.0017, clearly showing the benefit of the first-person perspective. One example of the results from using first-person images for the condition is shown in Fig. 8.
It seems to be more difficult to directly predict the joints instead of recreating the first-person image, despite both requiring a spatial understanding through the third-person image. One possible reason for that is the redundancy in the arm [28, 29, 30]. Allowing for many poses that look similar while having very different joint configurations. Fig. 8 shows one particular interesting result demonstrating this, where the first-person perspective looks very similar, while the actual pose looks quite different. With so many very similar-looking poses, the model tends to predict values closer to the mean.
@@ V Conclusion
We observed a difference between quantitative analysis and qualitative analysis regarding the effect of iterative denoising steps. The quantitative analysis did not show clear improvements from using a higher number of noise steps, while the qualitative reveals a clear effect, especially the later denoising steps. The benefit of the later denoising steps was particularly high for the more complex datasets, showing how diffusion models have unique qualities that can be helpful when handling complex data. The difference between the quantitative and qualitative results also highlights a need for further image comparison metrics that can evaluate higher levels of abstractions beyond pixel-wise similarities.
The benefits of generating first-person images were shown through their use in generating joint values. A significant difference in performance was observed between using third-person perspective images and first-person perspective images. Based on this, we assume that this way of generating first-person perspectives can be useful for other approaches that work on imitation, be it as an additional modality to an existing third-person perspective or as a way of generating more training data. Since higher amounts of data did not have a positive effect on the results, we assume that having a noisy version of the desired output available during training can speed up the time it takes the model to learn consistent patterns across the datasets.
Throughout this paper, we have demonstrated that our approach effectively learns spatial relations and can resize and rotate objects to adjust them to a new perspective. This enables our approach to generate first-person perspective images from third-person perspective images, that are nearly indistinguishable from actual first-person perspective images, and surpass established approaches such as pix2pix or CycleGAN. The architecture has shown promising results across three different datasets, each with unique properties.